# OBELiX: A Curated Dataset of Crystal Structures and Experimentally Measured Ionic Conductivities for Lithium Solid-State Electrolytes

### v1

@@ Abstract
Solid-state electrolyte batteries are expected to replace liquid electrolyte lithium-ion batteries in the near future thanks to their higher theoretical energy density and improved safety. However, their adoption is currently hindered by their lower effective ionic conductivity, a quantity that governs charge and discharge rates. Identifying highly ion-conductive materials using conventional theoretical calculations and experimental validation is both time-consuming and resource-intensive. While machine learning holds the promise to expedite this process, relevant ionic conductivity and structural data is scarce. Here, we present OBELiX, a domain-expert-curated database of $\sim$600 synthesized solid electrolyte materials and their experimentally measured room temperature ionic conductivities gathered from literature. Each material is described by their measured composition, space group and lattice parameters. A full-crystal description in the form of a crystallographic information file (CIF) is provided for $\sim$320 structures for which atomic positions were available. We discuss various statistics and features of the dataset and provide training and testing splits that avoid data leakage. Finally, we benchmark seven existing ML models on the task of predicting ionic conductivity and discuss their performance. The goal of this work is to facilitate the use of machine learning for solid-state electrolyte materials discovery.
Solid-state, batteries, electrolytes, ionic conductivity, energy storage, crystal, inorganics
@@ 1 Introduction
Lithium-ion batteries (LIBs) used in most consumer electronics and electric vehicles have seen immense progress in terms of energy density, power density, safety and durability. However, their performance is reaching a plateau. Solid-state batteries are regarded as the next generation of batteries that may allow significant improvement over these characteristics (Janek & Zeier, 2016, 2023). The key difference between these two technologies is their electrolyte, the medium which allows the transport of ions during charge and discharge. A solid-state electrolyte (SSE)—as opposed to a liquid electrolyte in LIBs—permits new design choices that ultimately lead to better battery properties (Betz et al., 2019), let alone the fact that solid electrolytes are not flammable unlike their liquid counterparts.
SSEs have a long research history, starting from Faraday’s first discovery of fast ion transport in $\beta$-$\text{PbF}_{2}$ and $\text{Ag}_{2}\text{S}$ about 200 years ago (Funke, 2013), until the relatively recent development of Li-based ionic conductors, such as $\text{Li}_{10}\text{GeP}_{2}\text{S}_{12}$ (Kamaya et al., 2011).
@@
Ionic conductivity ($\sigma$), expressed in siemens per centimeter (S/cm), measures how easily ions can move through a medium or material. Ideal SSEs, also called “superionic” or “fast-ionic” conductors, are electrolytes that exhibit ionic conductivity comparable to those observed in liquid electrolytes and molten solids ($>$ 1 mS/cm). Only a limited number of room temperature ideal SSEs are known thus far within a small number of classes of materials such as LISICON (e.g., $\text{Li}_{14}\text{ZnGe}_{4}\text{O}_{16}$), NASICON (e.g., $\text{Li}_{1.3}\text{Al}_{0.3}\text{Ti}_{1.7}(\text{PO}_{4})_{3})$, garnet (e.g., $\text{Li}_{7}\text{Li}_{3}\text{Zr}_{2}\text{O}_{12}$), perovskite (e.g., $\text{Li}_{0.5}\text{La}_{0.5}\text{TiO}_{3}$), and argyrodite (e.g., $\text{Li}_{6}\text{PS}_{5}\text{Cl}$) (Janek & Zeier, 2023).
@@
Until now, the discovery of novel SSEs has largely relied on an incremental, experimental approach which consists, for example, of substituting atoms and elements in known compounds. This has allowed the discovery of some highly ion-conductive materials, but greatly limits the search space given that the experimental synthesis and characterization of a new, stable, inorganic solid-state electrolyte is a difficult and costly process that can take months to years (Zhao et al., 2022).
@@
Table 1: Comparison of our dataset (OBELiX) with existing ones based on key features and labels. For features, the numbers represent the number of entries with that feature that are labeled with at least one experimental or computational ion transport property (not necessarily ionic conductivity). Numbers in parentheses represent proprietary data.
@@
Sendek et al. (2017) 0 0 317 317 317 317
Jalem et al. (2018) 0 0 318 318 318 318
He et al. (2020) (SPSE) n.a. 12,000 75 (7,678) 75 (7,678) 75 (7,678) 75 (7,678)
Hargreaves et al. (2023) (LiIon) 465 820 820 0 0 0
Laskowski et al. (2023) 1346 1346 1346 0 (344) 0 (344) 0 (344)
Shon & Min (2023) n.a. 4032 4032 0 0 0
Yang et al. (2024) (DDSE) 1939 2448 2448 0 0 0
@@
Computational discovery, on the other hand, requires time-consuming atomistic simulations, such as ab initio molecular dynamics (AIMD), to accurately capture the complex relationship between ionic conductivity and the material’s structure and composition (Ceder et al., 2018; Qi et al., 2021; Bielefeld et al., 2020). These calculations can take from several hours to a few days for a single ionic conductivity and their parameters are often materials specific. Therefore, they are not well suited for large-scale explorations of hypothetical materials.
Machine learning (ML) has the potential to greatly accelerate the discovery of novel SSEs. Naturally, it can be used to predict ionic conductivity directly using, for example, graph neural networks (GNNs), which have been used extensively and successfully in materials science (Schmidt et al., 2019; Butler et al., 2018). Machine-learned force fields or interatomic potentials (MLFF or MLIP) can also be used to obtain ionic conductivity through molecular dynamics in the “classical” way while using significantly less resources (Wines & Choudhary, 2024). Finally, generative frameworks can accelerate dynamics simulations (Nam et al., 2024) and, provided that good ionic conductivity models are developed, there exists a wide range of frameworks that could generate new materials conditioned on that property (Hernandez-Garcia et al., 2023; Zhu et al., 2024; Zeni et al., 2023; Merchant et al., 2023). However, the main obstacle to the development and validation of these models—and to some extent theoretical models—is the scarcity of relevant experimental ionic conductivity and structural datasets. Indeed, as detailed in the next section, the few datasets that exist contain partial material descriptions and ionic conductivity measurements at various temperatures. To the best of our knowledge there does not exist another open access dataset of experimental room temperature ionic conductivities with corresponding full crystal descriptions.
In this work, we assembled OBELiX (Open solid Battery Electrolytes with Li: an eXperimental dataset), a curated database of 599 synthesized solid electrolyte materials and their experimentally measured room temperature ionic conductivity along with descriptors of their space group, lattice parameters, and chemical composition [^1] OBELiX is available here: github.com/NRC-Mila/OBELiX. The database is analyzed in terms of the distribution of ionic conductivity, space groups, elements, and repeated compositions. We also propose a training and testing split that avoids data leakage between similar entries while balancing distributions of properties across splits. We use this split to benchmark the performance of 7 machine learning models at directly predicting room temperature ionic conductivity ($\sigma_{\text{RT}}$). Our database and benchmarks aim to significantly accelerate the ML-assisted discovery of novel SSEs with fast ion diffusion.
2 Related work
Crystal structure databases such as the Materials Project (Jain et al., 2013) or the Inorganic Crystal Structure Database (ICSD) (Belsky et al., 2002; Hellenbrandt, 2004) contain large amounts of potential candidates for solid-state electrolytes. For example, Sendek et al. (2017) screened more than 12,000 Li-containing crystals for Li-ion SSEs using multiple criteria, thereby identifying 317 candidates, among which 21 crystals that showed promise as SSEs were selected from an ML-guided model. The ionic conductivity of these 21 structures was computed theoretically.
Jalem et al. (2018) annotated 318 compounds by calculating ion migration energy barriers ($E_{b}$), a less accurate but computationally lighter property that relates to ionic conductivity. Bayesian optimization was employed to screen candidate compounds with low $E_{b}$.
He et al. (2020) compiled a database of over 90,000 crystal structures, including more than 7,000 structures with preliminary ion-transport data obtained through geometric analysis, and 12,000 activation energy values ($E_{b}$) calculated using the bond valence site energy method. Additionally, they manually extracted 75 CIF files from literature data. They employed empirical and geometrical methods to estimate the minimum energy paths of these structures and obtain $E_{b}$, but they did not predict $\sigma$.
In the last two years, four experimentally-derived datasets curated from a large collection of SSE literature were published.
@@ 2 Related work
Laskowski et al. (2023) gathered a dataset of 1346 entries with compositions, space group, and corresponding $\sigma_{\text{RT}}$, with a subset of 344 compounds whose structures are manually matched with an ICSD ID.
Finally, Shon & Min (2023) used text mining to extract more than 4000 ionic conductivity measurements from 1457 papers. Each ionic conductivity measurement is associated with a composition and about 350 are also associated with a “structure type”. Measurement temperature is not specified and compositions are not always fully described.
A recent study by Yang et al. (2024) introduced the Dynamic Database of Solid-State Electrolyte (DDSE) to facilitate the exploration of structure-performance relationships and accelerate the discovery of high-performance solid-state electrolytes (SSEs). The database contains performance data for 2448 materials (at time of writing), including ionic conductivity obtained from experimental reports, across a broad temperature range (132.40–1261.60 K). By employing data mining and machine learning, the DDSE provides tools to evaluate and predict ionic conductivity for new materials.
These recent reports greatly increased the amount of readily available experimental ionic conductivity data. However, they contain limited structural information: the databases by Shon & Min (2023) and Yang et al. (2024) contain only a qualitative structure description for some materials, the LiIon dataset only includes the structural family and the dataset by Laskowski et al. (2023) is limited to space group information. Although the full crystallographic information of the 344 compounds of the Laskowski dataset for which the ICSD ID is provided could be retrieved, the proprietary ICSD is not available to most researchers in the ML community. Table 1 summarizes the differences in terms of available features across the databases discussed above.
@@
Figure 2: a) Distributions of ionic-conductivity values for the training and testing sets along with proportions of crystal families and space groups. Only space groups that represent more than 1% of the sets are labeled. b) Venn diagram showing shared entries between our dataset and others c) Proportion of entries that contain each element in the periodic table. Elements that are not present in the dataset are shaded. Generated with pymatviz (Riebesell et al., 2022).
@@ 3.1 Background
While the combination of the composition, lattice parameters and space group is often sufficient to qualify materials, they do not fully describe the crystal structure because in general they do not specify the positions of each atom. Some experimental papers perform an analysis (Rietveld refinement) of the X-ray powder diffraction pattern of their materials to estimate atomic positions. This information is necessary to perform molecular dynamics simulations, for example.
In contrast to theory-based data found in the Materials Project, for example, experimental compositions often feature fractional numbers (real numbers rather than integers) resulting from partially vacant sites or disorder associated with partial cation substitution. Consider, for example, composition K0.1Li0.9SbO3. At a specific location in the crystal (a site) there is a 90% probability of finding a lithium (Li) atom and a 10% probability of finding a potassium (K) atom. Site occupancy does not need to add up to one, and sites are often partially empty. Such partial occupancy is ubiquitously observed in Li-ion SSEs and it plays a crucial role in creating diffusion pathways. For example, the $\sigma_{\text{RT}}$ of tetragonal $\text{Li}_{7}\text{Li}_{3}\text{Zr}_{2}\text{O}_{12}$ with a space group of I41/acd (no. 142) is two orders of magnitude smaller than that of the same garnet framework of cubic $\text{Li}_{7}\text{Li}_{3}\text{Zr}_{2}\text{O}_{12}$ with Ia-3d (no. 230) (see Figure 1a). In this case, the disordering and partial occupation of Li (at the 96h site) promotes the Li-ion conduction. In the halide structure of $\text{Li}_{3}\text{InCl}_{6}$ (Figure 1b)), the substitution of one Li+ with the In3+ cation introduces two intrinsic vacancies, to which are attributed the high $\sigma_{\text{RT}}$ of that material. In sum, in order to screen SSEs with high $\sigma_{\text{RT}}$, it is highly desirable to include partial occupancy as a key feature of the materials.
@@ 3.2 Data Collection
Ionic conductivity is usually reported as a property of the materials in the powder form which, includes the effect of defects and grain boundaries. It is referred to as “total” ionic conductivity. The ionic conductivity of individual grain is sometimes reported as the “bulk” ionic conductivity. When both were available we recorded both. This is relevant because the total ionic conductivity of materials not only depends on their crystal structure but also on factors such as the size of particles.
@@
Because the ICSD is not an open database, we are not able to publish 292 of the CIF files and can only link our entries to their corresponding ICSD ID. However, to reach a broader audience, in agreement with the ICSD, we openly publish a set of 292 CIF files for which a normally distributed random noise with standard deviation 0.01 ($\epsilon\sim N(0,0.01)$) in fractional coordinates was added to the original atomic positions. This noise was added while making sure that the full symmetry of the crystal was preserved. We measured the effect of noise on model performance (see section 4) and found that it made little to no difference (see the SI for more details).
@@ 3.3 Data Splits
We hope that by building the train and test splits carefully the performance of models on new data can be assessed more meaningfully.
@@ 4 Benchmarks
We note that experimental data intrinsically embeds errors and uncertainty associated not only with various sources of measurement techniques but also with data extraction from figures and inconsistent labelling (e.g., bulk, grain boundary, or total ionic conductivity are often indistinguishably reported). Before assessing the performance of predictive models it makes sense to quantify the uncertainty (“performance”) of experimental data acquisition. Thankfully, our dataset contains 48 sets of compositions and space groups that have multiple entries, spanning a total of 122 entries. These entries and their corresponding ionic conductivities are plotted in Figure 3. The color represents the maximum difference in lattice parameters between any two entries of a same set. The maximum difference is of only 1.2% for all sets, which gives us confidence that grouped materials are in fact the same. This means that these materials were synthesized and their ionic conductivity measured two or more times, most likely by different researchers. This represents a unique opportunity to quantify experimental uncertainty and reproducibility. The inset of Figure 3 shows the distribution of log ionic conductivities with respect to the mean of each set of repeated materials. The root mean squared deviation from the set averages is of 0.63 log(S/cm) and the mean absolute deviation from the set medians is of 0.41 log(S/cm). The latter can be compared to the model’s mean absolute error when predicting log ionic conductivity and represents its lower bound. Therefore, any model that would be reported as having lower MAE that that value would most likely be over-trained.

### v2

@@ Abstract
Solid-state electrolyte batteries are expected to replace liquid electrolyte lithium-ion batteries in the near future thanks to their higher theoretical energy density and improved safety. However, their adoption is currently hindered by their lower effective ionic conductivity, a quantity that governs charge and discharge rates. Identifying highly ion-conductive materials using conventional theoretical calculations and experimental validation is both time-consuming and resource-intensive. While machine learning holds the promise to expedite this process, relevant ionic conductivity and structural data is scarce. Here, we present OBELiX, a database of $\sim$600 synthesized solid electrolyte materials and their experimentally measured room temperature ionic conductivities gathered from literature and curated by domain experts. Each material is described by their measured composition, space group and lattice parameters. A full-crystal description in the form of a crystallographic information file (CIF) is provided for $\sim$320 structures for which atomic positions were available. We discuss various statistics and features of the dataset and provide training and testing splits carefully designed to avoid data leakage. Finally, we benchmark seven existing ML models on the task of predicting ionic conductivity and discuss their performance. The goal of this work is to facilitate the use of machine learning for solid-state electrolyte materials discovery.
@@ 1 Introduction
Lithium-ion batteries (LIBs) used in most consumer electronics and electric vehicles have seen immense progress in terms of energy density, power density, safety and durability. However, their performance is reaching a plateau. Solid-state batteries are regarded as the next generation of batteries that may allow significant improvement over these characteristics [15, 16]. The key difference between these two technologies is their electrolyte, the medium which allows the transport of ions during charge and discharge. A solid-state electrolyte (SSE)—as opposed to a liquid electrolyte in LIBs—permits new design choices that ultimately lead to better battery properties [3], let alone the fact that they are not flammable unlike their liquid counterparts.
@@
Ionic conductivity ($\sigma$), expressed in siemens per centimeter (S/cm), measures how easily ions can move through a medium or material. Ideal SSEs, also called “superionic” or “fast-ionic” conductors, are electrolytes that exhibit ionic conductivity comparable to those observed in liquid electrolytes and molten solids (>> 1 mS/cm). Only a handful of room temperature ideal SSEs are known thus far within a small number of classes of materials: LISICON (e.g., $\text{Li}_{14}\text{ZnGe}_{4}\text{O}_{16}$), NASICON (e.g., $\text{Li}_{1.3}\text{Al}_{0.3}\text{Ti}_{1.7}(\text{PO}_{4})_{3})$, garnet (e.g., $\text{Li}_{7}\text{Li}_{3}\text{Zr}_{2}\text{O}_{12}$), perovskite (e.g., $\text{Li}_{0.5}\text{La}_{0.5}\text{TiO}_{3}$), and argyrodite (e.g., $\text{Li}_{6}\text{PS}_{5}\text{Cl}$) [16].
@@
Until now, the discovery of novel SSEs has largely relied on an incremental, experimental approach which consists, for example, of substituting atoms and elements in known compounds. This has allowed the discovery of some highly ion-conductive materials, but greatly limits the search space given that the experimental synthesis and characterization of a new, stable, inorganic solid-state electrolyte is a difficult and costly process that can take months to years [35].
Computational discovery, on the other hand, requires time-consuming atomistic simulations, such as ab initio molecular dynamics (AIMD) which is based on density functional theory (DFT), to accurately capture the complex relationship between ionic conductivity and the material’s structure and composition [6, 22, 4]. These calculations can take from several hours to a few days for a single ionic conductivity and their parameters are often materials specific. Therefore, they are not well suited for large-scale explorations of hypothetical materials.
Machine learning (ML) has the potential to greatly accelerate the discovery of novel SSEs. Naturally, it can be used to predict ionic conductivity directly using, for example, graph neural networks (GNNs), which have been used extensively and successfully in materials science [24, 5]. Machine-learned force fields or interatomic potentials (MLFF or MLIP) can also be used to obtain ionic conductivity through molecular dynamics in the “classical” way while using significantly less resources [31]. Finally, generative frameworks can accelerate dynamics simulations [21] and, provided that good ionic conductivity models are developed, there exists a wide range of frameworks that could generate new materials conditioned on that property [12, 36, 34, 20]. However, the main obstacle to the development and validation of these models—and to some extent theoretical models—is the scarcity of relevant experimental ionic conductivity and structural datasets. Indeed, as detailed in the next section, the few datasets that exist contain partial material descriptions and ionic conductivity measurements at various or unspecified temperatures. To the best of our knowledge there does not exist another open access dataset of experimental room temperature ionic conductivities with corresponding full crystal descriptions.
In this work, we assembled OBELiX (Open solid Battery Electrolytes with Li: an eXperimental dataset), a curated database of 599 synthesized solid electrolyte materials and their experimentally measured room temperature ionic conductivity along with descriptors of their space group, lattice parameters, and chemical composition [^1] OBELiX is available here: github.com/NRC-Mila/OBELiX. The database is analyzed in terms of the distribution of ionic conductivity, space groups, elements, and repeated compositions. We also propose a training and testing split that avoids data leakage between similar entries while balancing distributions of properties across splits. We use this split to benchmark the performance of 7 machine learning models at directly predicting room temperature ionic conductivity ($\sigma_{\text{RT}}$).
We believe that this dataset and benchmark can significantly spur the use of ML for the discovery of novel solid-state battery materials. It may be small but it is important to realize that the database represents a large fraction of all materials whose ionic conductivity has been characterized experimentally. Importantly, this database has been carefully curated by domain experts and formatted by machine learning scientists to facilitate its use by this community. Finally, we believe that this benchmark can encourage novel machine learning research tailored to low-data regimes.
2 Related work
@@
Table 1: Comparison of our dataset (OBELiX) with existing ones based on key features and labels. For features, the numbers represent the number of entries with that feature that are labeled with at least one experimental or computational ion transport property (not necessarily ionic conductivity). Numbers in parentheses represent proprietary or private data.
@@
Sendek et al. 0 0 317 317 317 317
Jalem et al. 0 0 318 318 318 318
He et al. (SPSE) 0 0 75 (12k) 75 (12k) 75 (12k) 75 (12k)
Hargreaves et al.(LiIon) 465 820 820 0 0 0
Laskowski et al. 1346 1346 1346 0 (344) 0 (344) 0 (344)
Shon and Min n.a. 4032 4032 0 0 0
Yang et al. (DDSE) (1939) (2448) 2448 0 0 0
@@
Crystal structure databases such as the Materials Project [13] or the Inorganic Crystal Structure Database (ICSD) [2, 11] contain large amounts of potential candidates for solid-state electrolytes. For example, Sendek et al. [28] screened more than 12,000 Li-containing crystals for Li-ion SSEs using multiple criteria, thereby identifying 317 candidates, among which 21 crystals that showed promise as SSEs were selected from an ML-guided model. The ionic conductivity of these 21 structures was estimated theoretically. Jalem et al. [14] annotated 318 compounds by calculating ion migration energy barriers ($E_{b}$), a less accurate but computationally lighter property that relates to ionic conductivity. Bayesian optimization was employed to screen candidate compounds with low $E_{b}$. He et al. [10] compiled a database of over 90,000 crystal structures, including more than 7,000 structures with preliminary ion-transport data obtained through geometric analysis, and 12,000 activation energy values ($E_{b}$) calculated using the bond valence site energy method. Additionally, they manually extracted 75 CIF files from literature data. They employed empirical and geometrical methods to estimate the minimum energy paths of these structures and obtain $E_{b}$, but they did not predict $\sigma$.
@@
These recent reports greatly increased the amount of readily available experimental ionic conductivity data. However, they contain limited structural information: the databases by Shon and Min [29] and Yang et al. [33] contain only a qualitative structure description for some materials, the LiIon dataset only includes the structural family and the dataset by Laskowski et al. [17] is limited to space group information. Although the full crystallographic information of the 344 compounds of the Laskowski dataset for which the ICSD ID is provided could be retrieved, the proprietary ICSD is not available to most researchers in the ML community. Table 1 summarizes the differences in terms of available features across the databases discussed above.
@@ 3.1 Background
While the combination of the composition, lattice parameters and space group is often sufficient to qualify materials, they do not fully describe the crystal structure because in general they do not specify the positions of each atom. Some, but not all, experimental papers perform an additional analysis (Rietveld refinement) of the X-ray powder diffraction pattern to estimate atomic positions. Only in these cases is it possible to obtain a full description of the crystal including atomic positions which is necessary to build a crystal information file (CIF). This is why it is not possible to obtain CIF files for all entries in our dataset. The full crystal description including atomic positions is the information required to perform, for example, molecular dynamics simulations or density functional theory calculations.
In contrast to theory-based data found in the Materials Project, for example, experimental compositions often feature fractional numbers (real numbers rather than integers) resulting from partially vacant sites or disorder associated with partial cation substitution. Consider, for example, composition K0.1Li0.9SbO3. At a specific location in the crystal (a site) there is a 90% probability of finding a lithium (Li) atom and a 10% probability of finding a potassium (K) atom. Site occupancy does not need to add up to one since sites are often partially empty.
Such partial occupancy is ubiquitously observed in Li-ion SSEs [19] and it plays a crucial role in creating diffusion pathways. For example, the $\sigma_{\text{RT}}$ of tetragonal $\text{Li}_{7}\text{Li}_{3}\text{Zr}_{2}\text{O}_{12}$ with a space group of I41/acd (no. 142) is two orders of magnitude smaller than that of the same garnet framework of cubic $\text{Li}_{7}\text{Li}_{3}\text{Zr}_{2}\text{O}_{12}$ with Ia-3d (no. 230) (see Figure 1a). In this case, the disordering and partial occupation of Li (at the 96h site) promotes the Li-ion conduction. In the halide structure of $\text{Li}_{3}\text{InCl}_{6}$ (Figure 1b), the substitution of one Li+ with the In3+ cation introduces two intrinsic vacancies, to which is attributed the high $\sigma_{\text{RT}}$ of that material. In sum, in order to screen SSEs with high $\sigma_{\text{RT}}$, it is highly desirable to include partial occupancy as a key feature of the materials.
@@ 3.2 Data collection
Ionic conductivity is usually reported as a property of the materials in the powder form, which includes the effect of defects and grain boundaries. It is referred to as “total” ionic conductivity. The ionic conductivity of individual grain is sometimes reported as the “bulk” ionic conductivity. When both were available we recorded both. This is relevant because the total ionic conductivity of materials not only depends on their crystal structure but also on factors such as the size of particles.
@@
Because the ICSD is a proprietary database, we are not able to publish 292 of the CIF files and can only link our entries to their corresponding ICSD ID. However, to reach a broader audience, in agreement with the ICSD, we openly publish a set of 292 CIF files for which a normally distributed random noise with standard deviation 0.01 ($\epsilon\sim N(0,0.01)$) in fractional coordinates was added to the original atomic positions. This noise was added while making sure that the full symmetry of the crystal was preserved. We measured the effect of noise on model performance (see section 4) and found that it made little to no difference (see the SI for more details).
Figure 2: a) Distributions of ionic-conductivity values for the training and testing sets along with proportions of crystal families and space groups. Only space groups that represent more than 1% of the sets are labeled. b) Venn diagram showing how OBELiX entries are shared across the ICSD, Laskowski and LiIon datasets. There are 2 OBELiX entries that are not part of any of the three datasets. c) Proportion of entries that contain each element in the periodic table. Elements that are not present in the dataset are shaded. Generated with pymatviz [23].
@@ 3.3 Data splits
About 75% (245/321) of the entries with atomic information have some level of partial occupancy (disorder). The proportion of partially occupied structures in each split was not controlled for explicitly, but it is similar in the test (53/67) and train (192/254) splits. For the rest of the entries, when atomic positions and occupations are unknown, it is not always possible to tell if a structure is disordered.
@@ 4 Benchmarks
We note that experimental data intrinsically embeds errors and uncertainty associated not only with various sources of measurement techniques but also with data extraction from figures and inconsistent labeling (e.g., bulk, grain boundary, or total ionic conductivity are often indistinguishably reported). Before assessing the performance of predictive models it makes sense to quantify the uncertainty (“performance”) of experimental data acquisition. Thankfully, our dataset contains 48 sets of compositions and space groups that have multiple entries, spanning a total of 122 entries. These entries and their corresponding ionic conductivities are plotted in Figure 3. The color represents the maximum difference in lattice parameters between any two entries of a same set. The maximum difference is of only 1.2% for all sets, which gives us confidence that grouped materials are in fact the same. This means that these materials were synthesized and their ionic conductivity measured two or more times, most likely by different researchers. This represents a unique opportunity to quantify experimental uncertainty and reproducibility. The inset of Figure 3 shows the distribution of log ionic conductivities with respect to the mean of each set of repeated materials. The root mean squared deviation from the set averages of the log($\sigma_{RT}$) is of 0.63 and the mean absolute deviation from the set medians is of 0.41. The latter can be compared to the model’s mean absolute error when predicting log ionic conductivity and represents its lower bound. Therefore, any model that would be reported as having lower MAE than that value would most likely be over-trained.